40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jul 26 20:12:18 2010, kiwamu, Update, SUS, OSEMs on PRM DSC_2279.JPG
    Reply  Tue Jul 27 21:38:21 2010, kiwamu, Update, SUS, installation of in-vac optics DSC_2284.JPGDSC_2291.JPG
       Reply  Wed Jul 28 17:32:37 2010, kiwamu, Update, SUS, installation of in-vac optics 
          Reply  Thu Jul 29 10:39:43 2010, kiwamu, Update, SUS, Re: installation of in-vac optics 
Message ID: 3313     Entry time: Wed Jul 28 17:32:37 2010     In reply to: 3302     Reply to this: 3315
Author: kiwamu 
Type: Update 
Category: SUS 
Subject: installation of in-vac optics 

[Jenne, Koji and Kiwamu]

We have installed two steering mirrors and the green periscope.

Also we took the tip-tilt out from the chamber for characterization.

1. installed two steering mirrors

     - We installed IPPOSSM2 and IPPOSSM3.

    - Because IPPOSSM2 is a new 0 deg mirror so we newly engraved "Y1-LW1-2037-1064-0-AR" on the mount and deleted  the previous enegravement.

      For the 0 deg mirror itself, it had already been engraved by Koji. The wedge was horizontally aligned.

      Now IPPOSSM2 is off from the right place by 5 inch for convenience because it may touch our stomachs when we try to lean further into the chamber.

    - IPPOSSM3 is a 45 deg mirror which used to be in the chamber and had been already correctly engraved, so we didn't have to newly engrave on it. Now it's on the right place approximately.

 2. put three oplev mirrors

      - Two of them are approximately on the place, but one which is going to be on the center of the table is not on the place because there is a cable distributer sitting exactly there.

 3. installed the green periscope

      - Both the mirrors and the periscope were correctly engraved.

     - Now it's sitting on the right place approximately.

 4. removed the tip-tilt 

      -  This tip-tilt is now in the clean room and the mechanical mode will be characterized. 


 The next things to be done

* Cross-coupling evaluation of the PRM OSEMs

* Rearrangement of the cable distributer panel.

   - In order to do this we have to pull its cables which are attached to the stack.

* Installation of three green steering mirrors

   - All of them need to be engraved.

* Installation of two tip-tilt

  - One for the SRC and the other for the PRC

ELOG V3.1.3-